# Computational Bifurcation Analysis

### v1

@@ 4.2. A cautionary note and its resolution
Computational tools designed to account for symmetries, such as symcon [21] and sympercon [16, 22], avoid the need to analyze near-singular problems by modifying the defining systems to explicitly account for the desired symmetry constraints. For example, sympercon eliminates the consequences of non-uniqueness and discretization near symmetry-breaking bifurcations by replacing the periodic-orbit boundary-value problem on the interval $[0,T]$ with a boundary-value problem on a shorter interval, say $[0,T/\ell]$, with boundary conditions that impose the desired symmetry. This approach not only avoids singularities associated with symmetry-breaking bifurcations, but also saves computational effort. The required modifications to the defining system, however, depend on the symmetry type.
@@ 4.3. Tracking bifurcations
[…]
@@ 5. Concluding discussion
The discretization of periodic orbits with collocation is especially suitable for time profiles with strongly non-harmonic shape such as those shown in Figure 5(d) with sharp peaks. Some problem classes, such as vibration analysis for mechanical structures, consider scenarios with simple (e.g., harmonic) forcing but complex spatial shapes. For these problems a general-purpose continuation method for continuation of periodic responses benefits from the spectral convergence of harmonic projections, such that harmonic balance with very few modes results in sufficient accuracy [28, 29]. For example, the vibrations of a satellite structure with $37$ degrees of freedom, used as an illustration in [29] could be analysed using just $5$ harmonics (comparing to $9$ harmonics to confirm accuracy). The harmonic balance method is also suitable for cases when the right-hand side can only be evaluated with low accuracy but high sampling rate (e.g., through experimental measurements), because the projection on harmonics has an averaging effect on random disturbances [30]. For this reason, it has been combined with feedback control to track bifurcations and unstable periodic responses in vibration experiments [31, 32]. See [33] for a review of methods for bifurcation analysis in controlled experiments.
For high-dimensional problems arising from discretizations of PDEs in fluid dynamics time-stepper based methods, combined with matrix-free linear algebra for evaluation of Jacobian matrices and linear equation solving, are often the most accessible option even for finding equilibria if the time stepper is already available [34, 35]. The review [34] refers to several demonstrations of this approach, including bifurcation analysis of stability boundaries and unstable states of a full 3D model of the Atlantic Meridional Overturning Circulation [36]. As methods based on time steppers using matrix-free linear algebra only need the existence of an arbitrarily initializable time-stepping algorithm for the dynamical system they have also been applied as entirely “equation-free” computations in many-particle simulations in chemistry [37, 38] and in agent-based models [39].
A special case for continuation in PDEs is the tracking of patterns. The general-purpose package pde2path is discussed in Chapter LABEL:GohRademacherLloyd24 of this handbook [40], making bifurcation analysis possible for general classes of PDEs such as elliptic PDEs or reaction-diffusion equations in $1$ to $3$ space dimensions. In particular, steady or steadily moving PDE patterns (such as traveling or spiral waves) arise as relative equilibria in domains and equations with continuous symmetry. These patterns require a treatment that generalizes the approach Section 4.1 for discrete symmetries to continuous symmetries.

### v2

@@ 4.2. A cautionary note and its resolution
Computational tools designed to account for symmetries, such as symcon [21] and sympercon [16, 22], avoid the need to analyze near-singular problems by modifying the defining systems to explicitly account for the desired symmetry constraints. For example, sympercon eliminates the consequences of non-uniqueness and discretization near symmetry-breaking bifurcations by replacing the periodic-orbit boundary-value problem on the interval $[0,T]$ with a boundary-value problem on a shorter interval, say $0,T/\ell]$, with boundary conditions that impose the desired symmetry. This approach not only avoids singularities associated with symmetry-breaking bifurcations, but also saves computational effort. The required modifications to the defining system, however, depend on the symmetry type.
@@ 4.3. Tracking bifurcations
(20) $\begin{aligned} 0&=(\Pi\otimes I_{2}-E_{\ell}\otimes I_{2})(\sqrt{k}v_{1}-\mathrm{i}Jv_{1}),\end{aligned}$
@@ 5. Concluding discussion
The discretization of periodic orbits with collocation is especially suitable for time profiles with strongly non-harmonic shape such as those shown in Figure 5(d) with sharp peaks. Some problem classes, such as vibration analysis for mechanical structures, consider scenarios with simple (e.g., harmonic) forcing but complex spatial shapes. For these problems a general-purpose continuation method for continuation of periodic responses benefits from the spectral convergence of harmonic projections, such that harmonic balance with very few modes results in sufficient accuracy [28, 29]. For example, the vibrations of a satellite structure with $37$ degrees of freedom, used as an illustration in [29] could be analyzed using just $5$ harmonics (comparing to $9$ harmonics to confirm accuracy). The harmonic balance method is also suitable for cases where the right-hand side can only be evaluated with low accuracy but a high sampling rate (e.g., through experimental measurements), because the projection on harmonics has an averaging effect on random disturbances [30]. For this reason, it has been combined with feedback control to track bifurcations and unstable periodic responses in vibration experiments [31, 32]. See [33] for a review of methods for bifurcation analysis in controlled experiments.
For high-dimensional problems arising from discretizations of PDEs in fluid dynamics time-stepper based methods, combined with matrix-free linear algebra for evaluation of Jacobian matrices and linear equation solving, are often the most accessible option even for finding equilibria if the time stepper is already available [34, 35]. The review [34] refers to several demonstrations of this approach, including bifurcation analysis of stability boundaries and unstable states of a full 3D model of the Atlantic Meridional Overturning Circulation [36]. As methods based on time steppers using matrix-free linear algebra only need the existence of an arbitrarily initializable time-stepping algorithm for the dynamical system, they have also been applied as entirely “equation-free” computations in many-particle simulations in chemistry [37, 38] and in agent-based models [39].
A special case for continuation in PDEs is the tracking of patterns. The general-purpose package pde2path is discussed in Chapter LABEL:GohRademacherLloyd24 of this handbook [40], making bifurcation analysis possible for general classes of PDEs such as elliptic PDEs or reaction-diffusion equations in $1$ to $3$ space dimensions. In particular, steady or steadily moving PDE patterns (such as traveling or spiral waves) arise as relative equilibria in domains and equations with continuous symmetry. These patterns require a treatment that generalizes the approach in Section 4.1 for discrete symmetries to continuous symmetries.